• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name X. Zheng

    Log entry time 16:04:38 on March 3,2009

    Entry number 264344

    keyword=Shift Summary (Day)

    LHRS set for positron data-taking;
    Kin: H;
    momentum setting: 1.27 GeV/c;
    angle: 45 deg;
    beam energy:5.892 GeV/c.

    08:00 Shift starts with no beam.

    08:12 MCC says that from the beam point of view they are "ready" to send it but one of their Control computers for "STATUS" is down (this is the computer which tells where the beam is travelling), and can not send the beam; call for expert to bring the computer back.

    10:00 Going to controlled access.

    11:43 MCC said they are ready to send beam whenever we are ready.

    12:00 Jack is done with the Q3, back to beam permit.

    12:07 tuned beam.

    12:23 Harp scan result is 160x156um.

    12:26 turned on BB wire chamber HV. Checked rates with pol3He cell, 5uA, and raster sizes of 2x2.6, 4x5 and 6x6 (MCC readings).

    12:43 performed target NMR, 64%.

    12:50- started production run on kinematics H

    15:15 just noticed BB runs 1948-1950 don't have correct sync pattern. Killed CODA and tried rebooting ROC5,8,9,10. In a rush of rebooting we probablly rebooted something for the LHRS as well so we stopped the left DAQ too (got enough data for this kinematics already).

    The ROCs 8, 9 and 10 rebooted okay in the terminal but we couldn't connect to them in CODA.

    15:30 Bob fixed the BB CODA, meanwhile changing LHRS p0 to 1.13 GeV/c (Kine F).

    16:00 shift ends with taking production data at Kinematics F.

    run list:
    ------------------------------------
    Kinematics H (p_LHRS=1.270 GeV/c)
    Reference N2 cell run (122psi on display), raster 3x3mm (MCC 4x5):

    Run#, Runtime, # events. comments
    LHRS runs:
    20494, 12:47-12:59, 400K
    20495, 13:04-13:16, 380K
    BB run:
    1948, 12:47-13:16, 3.4M, sync problem

    Polarized 3He run, raster 3x3.5mm (MCC 5x6):
    Run#, Runtime, # events. comments
    LHRS runs:
    20496, 13:28-13:58, 400K
    20497, 13:58-15:00. 800K
    20498. 15:01-15:16. no end of run entry (some of the ROCs rebooted in the last 1-2 minutes).

    BB runs:
    1949. 13:28-14:17, 5M, sync problem
    1950. 14:20-15:12. 5M, sync problem
    1951. junk

    ------------------------------------
    Kinematics F (p_LHRS=1.130 GeV/c)
    20499. 15:41-

    BB runs:
    Polarized 3He run, raster 3x3.5mm (MCC 5x6):
    Run#, Runtime, # events. comments

    1952. 15:36-, sync problem